# Structural barriers to complete homogenization and wormholing in dissolving porous and fractured rocks

### v1

@@ 2.2 Reactive Transport
The two dimensionless groups we employ—the effective Damköhler number $\mathrm{Da}_{\mathrm{eff}}$ and the transport ratio GG—quantify the competition of advective, diffusive, and reactive time scales in both porous and fractured media. We define them with the total volumetric flow rate $Q$ rather than a mean velocity, because strong initial flow focusing in discrete fracture networks (DFNs) makes any “average” velocity unrepresentative. This volumetric definition also remains consistent for our pore–network simulations, which share identical width and an equal number of inlet pores, so $Q$ scales directly with the mean pore velocity.
A direct, one–to–one comparison of $\mathrm{Da}_{\mathrm{eff}}$ between the two systems is nevertheless impossible. In our models a cross-section perpendicular to the main flow contains only 200 conduits in the pore networks but roughly 5000 fracture elements in the DFNs; the vastly different hydraulic cross-sections imply different residence-time distributions even at the same global flow rate $Q$. Consequently, we treat $\mathrm{Da}_{\mathrm{eff}}$ as a system-specific control parameter, scanning it separately for pore and fracture networks and then comparing the resulting dissolution regimes rather than matching absolute values.
Simulation time is measured as the normalized dissolved volume, $T=\Delta V_{\textrm{tot}}/V_{0}$, where $V_{0}$ is the initial pore volume and $\Delta V_{\textrm{tot}}$ is the total dissolved volume in the system.
@@ 2.3 Flow Focusing Profile
The main tool that we use to determine the evolution of the system is the flow focusing profile [70]. We segment the medium into cross sections along the main flow direction, $x$, and in each of them we calculate the flow focusing index, $f_{50\%}$ [39], according to
@@
where $n_{50\%}$ is the smallest number of pore or fracture connections carrying 50% of the total flow through a given cross section of the medium, and $n_{\textrm{X}}$ is the total number of pore or fracture connections in that cross section. Calculating the index, $f_{50\%}$, for cross sections along the entire medium, we obtain a profile measuring flow focusing as a function of distance from the inlet at a given time.
@@ 3 Results
Each network type exhibits distinct dissolution behavior due to differences in structural organization and type of heterogeneity. For the regular pore network, all connections are of the same length, and only conduit-scale heterogeneity is present, in the form of the distribution of diameters. In the disordered pore network, apart from the conduit-scale heterogeneity, based on the same noise as in the regular network, there is also the segment-scale heterogeneity in the form of distribution of connection lengths and intersection positions. While in the case of porous media, the conduit- and segment-scale heterogeneity can be of the same order of magnitude, their couplings with dissolution are different, with the latter much more persistent. Finally, in the discrete fracture network, the nodes, connections, and apertures are distributed to approximate a real fracture network, introducing heterogeneity across all scales. Fracture families with various apertures introduce the conduit-scale heterogeneity, variance in fracture lengths generates the segment-scale heterogeneity (which in fractures can be much larger than that in apertures), and finally connectivity within the structure creates heterogeneity on the scale of the entire network, as sometimes only a few fractures can span the entire domain.
Figure 2: Evolution of the three network types in the wormholing regime. (a) Regular pore network. (b) Disordered pore network. For both pore lattices, edge width is proportional to the volumetric flow rate. (c) Discrete fracture network shown as a graph after a principal-component-analysis projection; edge width scales with flow, but only edges carrying more than 1% of the maximum flow rate in the system are plotted.
We simulate the three systems, choosing the dimensionless parameters $\textrm{Da}_{\textrm{eff}}$ and G to explore the standard dissolution regimes: uniform dissolution, channeling, and wormholing. Figure 2 presents the networks evolving in the wormholing regime, alongside the changes of the flow focusing profile. In pore networks, the dissolution regime becomes apparent almost immediately through the emergence of one or a few dominant channels. In the DFN the pattern is subtler—flow is further concentrated along pre-existing high-conductivity paths—but a newly developed high-permeability flow path can still be discerned. In each case the dissolution dynamics are captured by the flow focusing profile, which appears as a front of increased focusing that advances from the inlet toward the outlet; this increase is less pronounced in the DFN because strong focusing already exists at $T=0$.
@@ 3.1 Regular Pore Network
Figure 3: Evolution of the regular pore network and the flow focusing profile in three dissolution regimes: (a) uniform ($\textrm{Da}_{\textrm{eff}}=0.002$, $\textrm{G}=5$), (b) channeling ($\textrm{Da}_{\textrm{eff}}=0.02$, $\textrm{G}=5$), and (c) wormholing ($\textrm{Da}_{\textrm{eff}}=0.2$, $\textrm{G}=5$). Edge width is proportional to the volumetric flow rate.
@@
Figure 4: Evolution of the flow focusing profile for the regular pore network across the ($\textrm{Da}_{\textrm{eff}}$, G) parameter space. (a) Mean profile value at successive time snapshots (solid lines). (b) Same data, with ensemble variability: the shaded band around each curve denotes the mean $\pm$ 1 standard deviation across all realizations.
@@
We simulate the regular pore network for $\textrm{Da}_{\textrm{eff}}\in\{0.002,0.02,0.2\}$ and $\textrm{G}\in\{0.1,1,5\}$. The flow focusing profile is recorded at $T\in\{0,1,2,5,10\}$, giving a complete history of the dissolution process. Figure 3 presents the evolution of the network in the three dissolution regimes and Figure 4 shows the profile evolution for each $(\textrm{Da}_{\textrm{eff}},\textrm{G})$ pair: panel (a) displays the full range to reveal overall trends, whereas panel (b) focuses on the range of evolution of the profile, containing additional information on variability within the ensemble.
@@
In the uniform regime the flow focusing profile declines with time. The drop is steepest at high G: diffusion hindrance suppresses the growth of the widest channels, allowing narrower ones to dissolve faster so the network approaches complete homogeneity (profile values tend toward zero). At low G the values of the profile also decrease, but now because all channels enlarge at nearly the same rate, reducing their conductance contrasts. A small inlet–outlet asymmetry persists, indicating that the reactant penetration length—though large—is still finite relative to the system length.
@@
In the channeling regime, profile values increase in time along the entire system length. Initial main flow pathways widen almost uniformly from inlet to outlet, although the region near the inlet remains slightly more homogeneous, as more dissolution occurs there before the main flow pathway dominates the system. The effect of G is subtler. At low G the focusing weakens with depth, whereas at high G this decline nearly disappears, because diffusion hindrance allows more reactant to reach the deeper portions of the network. Conversely, at higher G the homogenization near the inlet is more pronounced, as dissolution in smaller channels is then faster.
@@
In the wormholing regime, a single highly conductive channel advances through the lattice, producing a sigmoidal profile: near-maximum values behind the tip and unchanged values ahead of it. The progress of the wormhole is faster for larger G—again, due to diffusion hindrance, more reactant is pushed toward the tip of the wormhole, instead of being used for widening the already existing parts of the channel.
@@
Figure 4b shows the variability of the flow focusing profile within the ensemble. The data indicate that, although different network realizations—which vary in their initial pore-diameter distributions—evolve at different rates, the variance becomes negligible by the end of the simulation. This convergence implies that the ultimate outcome—whether the degree of homogenization in the uniform regime or a characteristic level of channelization—does not depend on the specific initial diameter distribution.
@@ 3.2 Disordered Pore Network
Figure 5: Evolution of the disordered pore network and the flow focusing profile in three dissolution regimes: (a) uniform ($\textrm{Da}_{\textrm{eff}}=0.002$, $\textrm{G}=5$), (b) channeling ($\textrm{Da}_{\textrm{eff}}=0.02$, $\textrm{G}=5$), and (c) wormholing ($\textrm{Da}_{\textrm{eff}}=0.2$, $\textrm{G}=5$). Edge width is proportional to the volumetric flow rate.
@@
Figure 6: Evolution of the flow focusing profile for the disordered pore network across the ($\textrm{Da}_{\textrm{eff}}$, G) parameter space. (a) Mean profile value at successive time snapshots (solid lines). (b) Same data, with ensemble variability: the shaded band around each curve denotes the mean $\pm$ 1 standard deviation across all realizations.
@@
The results of dissolution for the disordered pore network are depicted in Figures 5 and  6. They are in agreement with our previous work [70]. We run the same set of simulations as for the regular lattice, using identical dimensionless parameters. We utilize the same diameter distributions, yet the initial flow focusing index in the ensemble is higher ($f_{50\%}\approx 0.5$ versus 0.4). This difference reflects the additional segment-scale heterogeneity introduced by non-uniform pore lengths and connectivity.
@@
The dissolution patterns evolve at different rates in the disordered and regular pore networks. In the channeling and wormholing regimes, the disordered network advances more slowly: at the same nondimensional time $T$, the channels in the disordered network extend noticeably less than those in the regular lattice. In the uniform regime, however, the disordered lattice dissolves at a comparable rate—and, at high G, even faster—than the regular network.
@@
Several factors contribute to the contrasting evolution of the two networks. Geometrically, the disordered lattice contains numerous connections perpendicular to the main pressure gradient. These cross-flow channels are nearly inactive in the uniform regime, where transverse pressure differences are negligible, but they start to carry substantial flow once wormholes form, because the wormhole tips impose strong lateral pressure gradients. This is increasing the reactant demand and slowing the growth. Second, pronounced aperture and length heterogeneity promotes branching, which further retards wormhole advance [10, 77]. When heterogeneity becomes extreme, however, stronger flow focusing can compensate for the branching penalty and accelerate growth again [10]. Because our disordered network displays significant heterogeneity at both conduit and segment scales, pinpointing its exact position on this spectrum—and how these competing effects balance—remains challenging.
@@
The two networks differ not only in how quickly they evolve but also in how completely they can homogenize. At low Damköhler numbers the disordered network still enters the uniform–dissolution regime, yet the resulting homogenization is weaker than in the regular network. Structural heterogeneity—independent of diameter variations—persists, and even at high G the flow–focusing index plateaus at $f_{50\%}\approx 0.25$. This saturation indicates an intrinsic limit to homogenization in the disordered medium. The snapshots in Fig. 5 change only marginally over time, confirming that much of the initial flow distribution is related to the disordered geometry of the network. This finding is crucial for studies aiming at complete homogenization: although uniform: although uniform dissolution can eliminate all diameter-based heterogeneity in regular lattices [63, 15], it cannot remove heterogeneity rooted in the network topology. Because natural rocks are never perfectly regular—both pore diameters and lengths vary—results derived from regular pore networks cannot be generalized directly to real systems.
@@
At intermediate $\textrm{Da}_{\textrm{eff}}$ the disordered network behaves differently from the regular lattice. Its extra heterogeneity—especially the variation in pore lengths—impedes the formation of dominant flow pathways spanning the entire system. Figure 5 shows that preferential channels exist from the outset, but they are far less pronounced than in the regular network. This is because the pore-diameter field is spatially correlated and thus encourages channel formation. Clusters of large-diameter pores create contiguous low-resistance corridors, so flow concentrates along them and dissolution amplifies the contrast, sharpening the channel. By contrast, the uncorrelated distribution of pore lengths work against sustained pathway development. As a result, the regime depends strongly on G. At small G, the network displays the features of the wormholing regime: a sigmoidal flow focusing profile and a dominant pathway that grows gradually from inlet to outlet. At higher G, the profile resembles that of the regular network, but with more initial homogenization. Because channel growth is slower, dissolution has more time to smooth out inlet-region heterogeneity before a single path takes over.
@@
At high $\textrm{Da}_{\textrm{eff}}$ both networks consistently develop wormholes and their growth accelerates with increasing $G$. The key difference lies in how strongly the wormholes compete. In the regular lattice, screening is stronger, so one wormhole soon outcompetes its neighbors and captures nearly the entire discharge. Once this conduit monopolizes the flow, the high velocity within it increases the reactant penetration length, broadening the transition zone in the flow focusing profile. In the disordered network, weaker screening allows several channels to share the flux; the dominant wormhole advances more slowly, and the focusing profile remains steeper.
@@
The ensemble variability shown in Figure 6b is of the same order of magnitude as in the regular network, but—because it now reflects spatially correlated diameter heterogeneity and the randomness of node positions—the profile is less smooth. How this variability evolves depends on the dissolution regime. In the uniform regime it decreases as conduit-scale heterogeneity is removed. In the channeling regime it first increases and then declines, indicating different evolution rates for individual realizations that eventually converge to a common outcome. In the wormholing regime variable growth rates are also evident; however, because the simulation ends before wormhole breakthrough, the expected late-time reduction in variability is not yet observed.

### v2

@@ 2.2 Reactive Transport
The two dimensionless groups we employ—the effective Damköhler number $\mathrm{Da}_{\mathrm{eff}}$ and the transport ratio G—quantify the competition of advective, diffusive, and reactive time scales in both porous and fractured media. We define them with the total volumetric flow rate $Q$ rather than a mean velocity, because strong initial flow focusing in discrete fracture networks (DFNs) makes any “average” velocity unrepresentative. This volumetric definition also remains consistent for our pore network simulations, which share identical width and an equal number of inlet pores, so $Q$ scales directly with the mean pore velocity.
A direct, one–to–one comparison of $\mathrm{Da}_{\mathrm{eff}}$ between the two systems is nevertheless impossible. In our models, a cross-section perpendicular to the main flow contains only 200 conduits in the pore networks but roughly 5000 fracture elements in the DFNs; the vastly different hydraulic cross-sections imply different residence-time distributions even at the same global flow rate $Q$. Consequently, we treat $\mathrm{Da}_{\mathrm{eff}}$ as a system-specific control parameter, scanning it separately for pore and fracture networks and then comparing the resulting dissolution regimes rather than matching absolute values.
Simulation time is expressed as a dimensionless dissolved volume,
$T=\frac{V_{\textrm{diss}}}{V^{\textrm{p}}_{0}},$ (9)
where $V^{\textrm{p}}_{0}$ is the initial total pore (or fracture) volume and $V_{\textrm{diss}}$ is the cumulative volume of solid that has dissolved. Thus, $T$ measures the total dissolved solid volume, normalized by the initial pore volume.
The above model is implemented in our in-house network simulation code, written in Python, as described in Szawełło, Szawełło [2025] where the source code and input files are openly available.
2.3 Model Assumptions and Limitations
The model relies on several simplifying assumptions and approximations, which we summarize here for clarity. First, we represent the porous medium as a network of cylindrical pipes and the fractured medium as a network of rectangular fracture segments. Both the geometry of each conduit and the intersections (which we treat as volumeless nodes) are simplified in this way. This idealization is standard in pore network and DFN models and is what allows us to extend modeling beyond the scale of individual pores while retaining analytical expressions for flow and transport within each conduit.
A second approximation concerns how dissolution modifies the conduit geometry. Because the reactant concentration decays along each conduit, the local dissolution rate is largest near the upstream end and smallest near the downstream end. In our model, we compute the total volume of mineral dissolved along the conduit and increase its diameter (or aperture) uniformly. This update is most accurate when the reactant penetration length,
$l_{\textrm{p}}=\frac{L}{\textrm{Da}_{\textrm{eff}}},$ (10)
is large compared with the average conduit length, $l_{\textrm{p}}\gg l_{0}$. The same scale enters the treatment of axial diffusion: neglecting diffusion along the flow direction is justified when the Péclet number based on $l_{\textrm{p}}$ is greater than one, so that advection dominates on the scale over which the reactant is consumed. This restriction means that the model is not intended to describe strongly diffusion-dominated regimes such as compact dissolution.
Additional assumptions are made regarding transport and chemical reactions. We assume a separation of time scales between flow and transport relaxation and mineral dissolution (acid capacity number $\gamma\ll 1$). We also assume perfect mixing of reactant at each intersection, which is a reasonable approximation for disordered networks but may overestimate mixing in simple regular lattices at high Péclet numbers, where streamline-based mixing models can be more appropriate. Sharma et al. [2023] Finally, the chemical kinetics are represented by a single-component linear rate law. This is a common approximation across a range of systems, but care is needed when applying it to systems with multiple reactive species and strong buffering, where equilibrium reactions can substantially modify the fluid chemistry (for example by buffering the $\textrm{H}^{+}$ concentration).
@@ 2.4 Flow Focusing Profile
The main tool that we use to determine the evolution of the system is the flow focusing profile. Szawełło et al. [2024] We segment the medium into cross sections along the main flow direction, $x$, and in each of them we calculate the flow focusing index, $f_{50\%}$Jang et al. [2011] according to
@@
where $n_{50\%}$ is the smallest number of conduits carrying 50% of the total flow through a given cross section of the medium, and $n_{\textrm{X}}$ is the total number of conduits in that cross section. Calculating the index, $f_{50\%}$, for cross sections along the entire medium, we obtain a profile measuring flow focusing as a function of distance from the inlet at a given time.
@@ 3 Results
Figure 2: Evolution of the three network types and the flow focusing profile in the wormholing regime. (a) Regular pore network ($\textrm{Da}_{\textrm{eff}}=0.2$, $\textrm{G}=5$). (b) Disordered pore network ($\textrm{Da}_{\textrm{eff}}=0.2$, $\textrm{G}=5$). (c) Discrete fracture network ($\textrm{Da}_{\textrm{eff}}=0.02$, $\textrm{G}=5$) shown as a graph after a principal-component-analysis projection. In each case, edge width is proportional to the volumetric flow rate, with the same proportionality constant used for a given network type. For the DFN, only edges carrying more than 1% of the maximum flow rate in the system are plotted. In the flow focusing plots, the initial profile is shown by the black line, and profiles at later times ($T=1.0,2.0,5.0,$ and $10.0$ for both pore networks; $T=0.1,0.2,0.5,$ and $1.0$ for the fracture network) are shown by colored lines.
Each network type exhibits distinct dissolution behavior due to differences in structural organization and type of heterogeneity. For the regular pore network, all conduits are of the same length, and only conduit-scale heterogeneity is present, in the form of the distribution of diameters. In the disordered pore network, apart from the conduit-scale heterogeneity, based on the same noise as in the regular network, there is also the segment-scale heterogeneity in the form of distribution of conduit lengths and intersection positions. While in the case of porous media, the conduit- and segment-scale heterogeneity can be of the same order of magnitude, their couplings with dissolution are different, with the latter much more persistent. Finally, in the discrete fracture network, the nodes, edges, and apertures are distributed to approximate a real fracture network, introducing heterogeneity across all scales. Fracture families with various apertures introduce the conduit-scale heterogeneity, variance in fracture lengths generates the segment-scale heterogeneity (which in fractures can be much larger than that in apertures), and finally connectivity within the structure creates heterogeneity on the scale of the entire network, as sometimes only a few fractures can span the entire domain.
We simulate the three systems, choosing the dimensionless parameters $\textrm{Da}_{\textrm{eff}}$ and G to explore the standard dissolution regimes: uniform dissolution, channeling, and wormholing. Fig. 2 presents the networks evolving in the wormholing regime, alongside the changes of the flow focusing profile. In pore networks, the dissolution regime becomes apparent almost immediately through the emergence of one or a few dominant channels. In the DFN the pattern is subtler—flow is further concentrated along pre-existing high-conductivity paths—but a newly developed high-permeability flow path can still be discerned. In each case the dissolution dynamics are captured by the flow focusing profile, which appears as a front of increased focusing that advances from the inlet toward the outlet; this increase is less pronounced in the DFN because strong focusing already exists at $T=0.0$.
@@ 3.1 Regular Pore Network
Figure 3: Evolution of the regular pore network and the flow focusing profile in three dissolution regimes: (a) uniform ($\textrm{Da}_{\textrm{eff}}=0.002$, $\textrm{G}=5$), (b) channeling ($\textrm{Da}_{\textrm{eff}}=0.02$, $\textrm{G}=5$), and (c) wormholing ($\textrm{Da}_{\textrm{eff}}=0.2$, $\textrm{G}=5$). Edge width is proportional to the volumetric flow rate, with the same proportionality constant used in all panels. The plots show the initial flow focusing profile (black line) and profiles at times $T=1.0,2.0,5.0,$ and $10.0$ (colored lines). The frame color denotes the dissolution regime: green for uniform, red for channeling, and blue for wormholing.
@@
Figure 4: Evolution of the flow focusing profile for the regular pore network across the ($\textrm{Da}_{\textrm{eff}}$, G) parameter space. (a) Mean profile value at successive time snapshots (solid lines). (b) Same data, with ensemble variability: the shaded band around each curve denotes the mean $\pm$ 1 standard deviation across all realizations. The plots show the initial flow focusing profile (black line) and profiles at times $T=1.0,2.0,5.0,$ and $10.0$ (colored lines). The frame color denotes the dissolution regime: green for uniform, red for channeling, and blue for wormholing.
@@
We simulate the regular pore network for $\textrm{Da}_{\textrm{eff}}\in\{0.002,0.02,0.2\}$ and $\textrm{G}\in\{0.1,1,5\}$. The flow focusing profile is recorded at $T\in\{0.0,1.0,2.0,5.0,10.0\}$, giving a complete history of the dissolution process. Fig. 3 presents the evolution of the network in the three dissolution regimes and Fig. 4 shows the profile evolution for each $(\textrm{Da}_{\textrm{eff}},\textrm{G})$ pair: panel (a) displays the full range to reveal overall trends, whereas panel (b) focuses on the range of evolution of the profile, containing additional information on variability within the ensemble.
@@
In the uniform regime (top rows of Fig. 4a,b), the flow focusing profile declines with time. The drop is steepest at high G: diffusion hindrance suppresses the growth of the widest channels, allowing narrower ones to dissolve faster so the network approaches complete homogeneity (profile values tend toward zero at $\textrm{G}=5$, as shown in the top-right panels of Fig. 4a,b). At low G the values of the profile also decrease, but now because all channels enlarge at nearly the same rate, reducing their conductance contrasts. A small inlet–outlet asymmetry persists, indicating that the reactant penetration length—though large—is still finite relative to the system length.
@@
In the channeling regime (middle rows of Fig. 4a,b), profile values increase in time along the entire system length. Initial main flow paths widen almost uniformly from inlet to outlet, although the region near the inlet remains slightly more homogeneous, as more dissolution occurs there before the main flow path dominates the system. The effect of G is subtler. At low G the focusing weakens with depth, whereas at high G this decline nearly disappears, because diffusion hindrance allows more reactant to reach the deeper portions of the network. Conversely, at higher G the homogenization near the inlet is more pronounced, as dissolution in smaller channels is then faster. This inlet-outlet asymmetry is clearly visible in the dissolution pattern in Fig. 3b.
@@
In the wormholing regime (bottom rows of Fig. 4a,b), a single highly conductive channel advances through the lattice, producing a sigmoidal profile: near-maximum values behind the tip and unchanged values ahead of it. The progress of the wormhole is faster for larger G—again, due to diffusion hindrance, more reactant is pushed toward the tip of the wormhole, instead of being used for widening the already existing parts of the channel.
@@
Fig. 4b shows the variability of the flow focusing profile within the ensemble. The data indicate that, although different network realizations—which vary in their initial pore-diameter distributions—evolve at different rates, the variance becomes negligible by the end of the simulation. This convergence implies that the ultimate outcome—whether the degree of homogenization in the uniform regime or a characteristic level of channelization—does not depend on the specific initial diameter distribution.
@@ 3.2 Disordered Pore Network
Figure 5: Evolution of the disordered pore network and the flow focusing profile in three dissolution regimes: (a) uniform ($\textrm{Da}_{\textrm{eff}}=0.002$, $\textrm{G}=5$), (b) channeling ($\textrm{Da}_{\textrm{eff}}=0.02$, $\textrm{G}=5$), and (c) wormholing ($\textrm{Da}_{\textrm{eff}}=0.2$, $\textrm{G}=5$). Edge width is proportional to the volumetric flow rate, with the same proportionality constant used in all panels. The plots show the initial flow focusing profile (black line) and profiles at times $T=1.0,2.0,5.0,$ and $10.0$ (colored lines). The frame color denotes the dissolution regime: green for uniform, red for channeling, and blue for wormholing.
@@
Figure 6: Evolution of the flow focusing profile for the disordered pore network across the ($\textrm{Da}_{\textrm{eff}}$, G) parameter space. (a) Mean profile value at successive time snapshots (solid lines). (b) Same data, with ensemble variability: the shaded band around each curve denotes the mean $\pm$ 1 standard deviation across all realizations. The plots show the initial flow focusing profile (black line) and profiles at times $T=1.0,2.0,5.0,$ and $10.0$ (colored lines). The frame color denotes the dissolution regime: green for uniform, red for channeling, and blue for wormholing.
@@
The results of dissolution for the disordered pore network are depicted in Figs. 5 and 6. They are in agreement with our previous work. Szawełło et al. [2024] We run the same set of simulations as for the regular lattice, using identical dimensionless parameters. We utilize the same diameter distributions, yet the initial flow focusing index in the ensemble is higher ($f_{50\%}\approx 0.5$ in Fig. 6 versus 0.4 in Fig. 4). This difference reflects the additional segment-scale heterogeneity introduced by non-uniform pore lengths.
@@
The dissolution patterns evolve at different rates in the disordered and regular pore networks, as shown by the patterns in Figs. 3 and 5. In the channeling and wormholing regimes, the disordered network advances more slowly: at the same nondimensional time $T$, the channels in the disordered network extend noticeably less than those in the regular lattice. In the uniform regime, however, the disordered lattice dissolves at a comparable rate—and, at high G, even faster—than the regular network.
@@
Several factors contribute to the contrasting evolution of the two networks. Geometrically, the disordered lattice contains numerous edges perpendicular to the main pressure gradient. These cross-flow channels are nearly inactive in the uniform regime, where transverse pressure differences are negligible, but they start to carry substantial flow once wormholes form, because the wormhole tips impose strong lateral pressure gradients. This is increasing the reactant demand and slowing the growth. Second, pronounced aperture and length heterogeneity promotes branching, which further retards wormhole advance. Cheung and Rajaram [2002], Upadhyay et al. [2015] When heterogeneity becomes extreme, however, stronger flow focusing can compensate for the branching penalty and accelerate growth again. Cheung and Rajaram [2002] Because our disordered network displays significant heterogeneity at both conduit and segment scales, pinpointing its exact position on this spectrum—and how these competing effects balance—remains challenging.
@@
The two networks differ not only in how quickly they evolve but also in how completely they can homogenize. At low Damköhler numbers (top rows of Fig. 6a,b), the disordered network still enters the uniform dissolution regime, yet the resulting homogenization is weaker than in the regular network. Structural heterogeneity—independent of diameter variations—persists, and even at high G the flow focusing index plateaus at $f_{50\%}\approx 0.25$, as shown in the top-right panels of Fig. 6a,b. This saturation indicates an intrinsic limit to homogenization in the disordered medium. The snapshots in Fig. 5a change only marginally over time, confirming that much of the initial flow distribution is related to the disordered geometry of the network. This finding is crucial for studies aiming at complete homogenization: although uniform dissolution can eliminate all diameter-based heterogeneity in regular lattices, Roded et al. [2020], Deng et al. [2025] it cannot remove heterogeneity rooted in the network topology. Because natural rocks are never perfectly regular—both pore diameters and lengths vary—results derived from regular pore networks cannot be generalized directly to real systems.
@@
At intermediate $\textrm{Da}_{\textrm{eff}}$ (middle rows of Fig. 6a,b), the disordered network behaves differently from the regular lattice. The additional heterogeneity—especially the variation in pore lengths—makes it harder for a single dominant flow path to form across the entire system. As shown in Fig. 5, preferential channels are present from the beginning, but they are much less pronounced than in the regular network. This reflects the different roles of the two types of heterogeneity. The pore-diameter field is spatially correlated, so clusters of large pores tend to line up and create continuous low-resistance corridors; flow concentrates along these corridors and dissolution then sharpens them into channels. The pore-length distribution, by contrast, is uncorrelated and therefore acts against the formation of such system-spanning paths. As a result, the dissolution regime depends strongly on G, as demonstrated by the behavior of the flow focusing profile in Fig. 6. At small G, the network displays the features of the wormholing regime: a sigmoidal profile and a dominant path that grows gradually from inlet to outlet. At higher G, the profile resembles that of the regular network, but with more initial homogenization. Because channel growth is slower, dissolution has more time to smooth out inlet-region heterogeneity before a single path takes over.
@@
At high $\textrm{Da}_{\textrm{eff}}$, both networks consistently develop wormholes and their growth accelerates with increasing G. The key difference lies in how strongly the wormholes compete. In the regular lattice, screening is stronger, so one wormhole soon outcompetes its neighbors and captures nearly the entire discharge. Once this conduit monopolizes the flow, the high velocity within it increases the reactant penetration length, broadening the transition zone in the flow focusing profile (bottom rows of Fig. 4a,b). In the disordered network, weaker screening allows several channels to share the flux; the dominant wormhole advances more slowly, and the focusing profile remains steeper (bottom rows of Fig. 6a,b).
@@
The ensemble variability shown in Fig. 6b is of the same order of magnitude as in the regular network, but—because it now reflects spatially correlated diameter heterogeneity and the randomness of node positions—the profile is less smooth. How this variability evolves depends on the dissolution regime. In the uniform regime it decreases as conduit-scale heterogeneity is removed. In the channeling regime it first increases and then declines, indicating different evolution rates for individual realizations that eventually converge to a common outcome. In the wormholing regime variable growth rates are also evident; however, because the simulation ends before wormhole breakthrough, the expected late-time reduction in variability is not yet observed.